OpenStudy (anonymous):

how can i know if {(x,y,z) E R^3 : z=xy} is a subspace? using : (for every constant) c,a E R(real number), (for every) w1, w2 E W : cw1 + aw2 E W

OpenStudy (unklerhaukus):

yeah

OpenStudy (anonymous):

yeah...? what yeah?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

frickin trolls

OpenStudy (watchmath):

It is not a subspace!

OpenStudy (anonymous):

how'd you find out using the condition i used

OpenStudy (watchmath):

(1,1,1) is belong to the set because 1=1*1. But 2(1,1,1)=(2,2,2) is not in the set since \( 2\neq 2\cdot 2\)

OpenStudy (anonymous):

ok... what about : z=x+y

OpenStudy (watchmath):

it is a subspace. You need to check it though :D

OpenStudy (anonymous):

if its (1,1,1) then it has to be 1=1+1, right? which is obviously wrong..

OpenStudy (watchmath):

if you want to say that it is not a subspace you can work with example. But if you want to show that it is a subspace you need to work in general

OpenStudy (anonymous):

i see...because i proved it IS a subspace, but when i think of it as (1,1,1) it doesn't make sense

OpenStudy (watchmath):

Let (a,b,c), (d,e,f) in the set. It means c=a+b and f=d+e. Let k,l be any scalar. Is k(a,b,c)+l(d,e,f) in the set? (how to check it? check it the third component is equal or not with the sum of the first two components)

OpenStudy (anonymous):

ka+kb=kc and ld+le=lf

OpenStudy (watchmath):

yes, but what do you need to show is that (ka+ld,kb+le,kc+lf) is in the set.

OpenStudy (anonymous):

ok so.. how

OpenStudy (watchmath):

what is the criterion if a vector belong to your set?

OpenStudy (anonymous):

uh?

OpenStudy (anonymous):

i don't understand what you mean

OpenStudy (anonymous):

z needs to equal x+y ..that?

OpenStudy (watchmath):

yes, exactly!

OpenStudy (watchmath):

but you need to read it as third component = first component + second component is that true for the vector (ka+ld,kb+le,kc+lf) ?

OpenStudy (anonymous):

no

OpenStudy (watchmath):

why not ?

OpenStudy (anonymous):

oh yes, it is

OpenStudy (watchmath):

why it is? :D

OpenStudy (anonymous):

first component was ka then ld, then kb and le, etc..

OpenStudy (watchmath):

you need to show that kc+lf = (ka+ld)+(kb+le)

OpenStudy (anonymous):

which is true..

OpenStudy (anonymous):

lets say: w1=(x1, y1, z1) and w2=(x2,y2,z2) then k(x1,y1,z1) + c(x2,y2,z2) = (kx1, ky1, kz1) + (cx2, cy2, cz2) = (kx1+cx2, ky1+cy2, kz1+kz2) so... (kx1+cx2) + (cy1+cy2) = kz1+kz2

OpenStudy (anonymous):

QED?

OpenStudy (watchmath):

almost you need to give an argument why (kx1+cx2) + (cy1+cy2) = kz1+kz2 is true

OpenStudy (anonymous):

hoooow xD

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

i kinda f'd up with the c and k

OpenStudy (anonymous):

k(x1,y1,z1) + c(x2,y2,z2) = (kx1, ky1, kz1) + (cx2, cy2, cz2) = (kx1+cx2, ky1+cy2, kz1+cz2) so... (kx1+cx2) + (cy1+cy2) = kz1+cz2

OpenStudy (anonymous):

thats corrected

OpenStudy (watchmath):

still a typo :). It should be (kx1+cx2) + (ky1+cy2) = kz1+cz2

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

anyway, so its: kx1+ky1= kz1 and cx2 + cy2= cz2

OpenStudy (anonymous):

that it?

OpenStudy (watchmath):

good! lets write all the argument neatly

OpenStudy (watchmath):

Let W={(x,y,z) : z=x+y}. Let : w1=(x1, y1, z1) and w2=(x2,y2,z2) are elements of W. Then z1=x1+y1 and z2=x2+y2. Now we want to show that k(x1,y1,z1)+c(x2,y2,z2) in W. Notice that k(x1,y1,z1)+c(x2,y2,z2)=(kx1+cx2,ky1+cy2,kz1+cz2). Now kz1+cz2=k(x1+y1)+c(x2+y2)=(kx1+cx2) +(ky1+cy2). Therefore k(x1,y1,z1)+c(x2,y2,z2) in W.

OpenStudy (anonymous):

i'm so smart..

OpenStudy (anonymous):

seriously though, thanks!

OpenStudy (anonymous):

the original question kinda got discarded, but its ok